Please show work for both questions5. The current in the 20 12 resistor in the circuit below is 1512 A. 1.5 A B. 5 A C. 10 A D. 15 A 30 A 3 20126. The current i, in the circuit below is A. 5 A B. - 5 A C. 3 A D. - 3 A 10V 222 +)3 is 3n3v,
Answer #1

Isu. 30A KCL > 30-inti Si= 30-10-30 KVL => -10(20) +i (15) +asip =0. =-sig+ (30-1)) #5 = 0. 5-sin-Asid= -30(85) > Poin = 30(66. rsis Ś Vo 3м - Applying mesh analysis we get 0 % -10+ is 2 =0 * sig = 5A: $-3i5 +218 +3 10 =0 => 3ig : 510 35 3(5)=sio opt